# Anomalous charge density wave in two-dimensional altermagnet WO

### v1

@@ Abstract
Recently, the study of novel physical properties arising from the combination of altermagnetism and other matter phases has attracted widespread attention, such as the integration of altermagnetism and topology. However, research on the combination of altermagnetism and charge density waves remains relatively sparse. In this letter, based on symmetry analysis and first-principles calculations, we demonstrate for the first time that altermagnetism and charge density waves can coexist in a two-dimensional material and predict monolayer WO to be such a material. Moreover, our calculations reveal that the altermagnetic order in monolayer WO stabilizes the charge density wave. Further, the charge density wave is not driven by Fermi-surface nesting but rather by strong electron–phonon coupling. More importantly, the charge density wave in monolayer WO leads to an anomalous transition from semimetal to metal. Therefore, we realize an anomalous charge density wave phase in altermagnetic WO. Considering the strong electron-phonon coupling and good metallic properties in the altermagnetic charge density wave state, our work may provide new insights into the realization of nontrivial altermagnetic superconductivity.
@@
Introduction. Altermagnetism, a new magnetic phase distinct from both ferromagnetism and antiferromagnetism, has been proposed theoretically and confirmed experimentally [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Altermagnetic materials exhibit momentum-space anisotropic spin splitting without spin-orbital coupling (SOC) and real-space antiparallel spin alignment with zero net total magnetic moment, thereby sharing the advantages of ferromagnetic and antiferromagnetic materials. Consequently, numerous exotic phenomena can emerge in altermagnetic materials, such as anisotropic spin splitting leading to spin currents [11, 12, 13, 14], giant and tunneling magnetoresistance [15, 16, 17], and piezomagnetic effects [18]. With SOC, due to the break of the time-reversal symmetry, altermagnetic materials can realize the anomalous Hall [19, 20, 21, 22], anomalous magneto-optical [23, 24], anomalous Nernst [25], anomalous thermal Hall [26] and crystal valley Hall [27] effects.
The combination of altermagnetism with other matter phases has also attracted widespread interest. The anisotropic spin-splitting altermagnetism with time-reversal symmetry merged with topology gives rise to new topological phases such as Chern insulators [21, 28], bipolarized Weyl semimetals [29], time-reversal-broken Dirac semimetals [30], higher-order topological insulators [31], and type-II quantum spin Hall insulators [32, 33]. Coupling time-reversal-broken altermagnetism with superconductivity can lead to nontrivial superconducting states [34, 35]. Moreover, the antiparallel spin alignment in altermagnets readily stabilizes insulating states with high Néel temperatures, so altermagnetic ferroelectricity is physically achievable and enables electrical control of magnetism [36, 37, 38, 39]. On the other hand, charge-density wave (CDW) can lower symmetry of materials [40, 41, 42, 43, 44], potentially transforming conventional collinear antiferromagnets into altermagnets. Thus, altermagnetism and CDW may coexist in a single material. Moreover, CDW in altermagnets have been investigated in three-dimensional materials [45, 46], but they remain unexplored in two-dimensional materials. CDW are known to emerge easily in two-dimensional materials [47], and altermagnetism has also been realized in two-dimensional materials [48, 49]. Therefore, realizing CDW phase in two-dimensional altermagnetic materials and exploring its novel physics is both important and urgent.
In this letter, based on symmetry analysis and first-principles calculations, we demonstrate for the first time that altermagnetism and CDW can coexist in a two-dimensional material, and predict monolayer WO to be such a material. Moreover, our calculations reveal that the altermagnetic order stabilizes the CDW in monolayer WO. Further, the CDW is not driven by Fermi-surface nesting but rather by strong electron–phonon coupling. Most importantly, the CDW lead to an anomalous transition from semimetal to metal in monolayer WO. Therefore, we realize an anomalous CDW phase in altermagnetic WO.
@@
Results and analysis. Recently, monolayer CrO and MoO have been predicted to be bipolarized Weyl semimetals with altermagnetism, and they will transition to the Chern insulator phase under strain [50, 21, 51]. Monolayer CrO is a perfect bipolarized Weyl semimetal, while the overlap of the valence and conduction bands in monolayer MoO leads to the presence of both electron and hole Fermi surfaces. Therefore, monolayer MoO exhibits better metallic properties than monolayer CrO. If W atoms replace Mo atoms, monolayer WO should have stronger metallic properties. Additionally, 5 transition metal oxides may lead to structural distortions [52], so monolayer WO provides the possibility for the coexistence of altermagnetism and CDW.
Similar to the monolayer CrO, the crystal structure of monolayer WO has the P4/mmm space group symmetry, with the corresponding point group being the D4h group. The unit cell of monolayer WO contains two W atoms and two O atoms (Fig. 1(a)), with an optimized lattice constant of 3.696 Å. To determine the magnetic structure of monolayer WO, we considered five magnetic structures, including one ferromagnetic and four antiferromagnetic configurations (Fig. 1(b)-(f)). Our calculations show that the AFM1 state remains the most stable as the correlation interaction U varies (Fig. 1(g)). For the AFM1 state, the two W atoms with opposite magnetic moments are respectively located at space-inversion invariant positions, giving WO the spin symmetry but breaking the spin symmetry . Due to the presence of O atoms, WO also breaks the spin symmetry . Here, C2 and represent the 180-degree rotation operation perpendicular to the spin in spin space and the fractional translation operation of in lattice space, respectively. In addition, the two W atoms with opposite magnetic moments can be connected by symmetry, which corresponds to spin symmetry . Therefore, monolayer WO is a -wave altermagnetic material, consistent with monolayer CrO and MoO.
@@
Next, we investigate the dynamical stability of monolayer WO based on altermagnetic structure. We calculated the phonon spectrum of WO, which is shown in Fig. 2(a). From Fig. 2(a), WO has two phonon modes with imaginary frequencies respective at the high-symmetry M and points, and the imaginary frequency at the M point is larger than that at the point. We speculate that the imaginary frequencies might be caused by a CDW. To verify this speculation, we first calculated the phonon spectrum of the supercell. Under the supercell, the M point of the primitive cell Brillouin zone (BZ) is folded to the point, as shown in Fig. 2(b). Therefore, the vibration mode with imaginary frequency at the M point of the primitive cell BZ is folded to the point. Then, we plotted these two vibration modes with imaginary frequency, which are shown in Fig. 2(c) and Fig. 2(d), respectively. From Fig. 2(c) and Fig. 2(d), these two vibration modes originate from the vibrations of W atoms. The vibration mode with imaginary frequency at the point is the relative vibration of W atoms along the z-direction, and the vibration mode with imaginary frequency at the M point is the in-plane relative vibration of W atoms. Based on these two vibration modes, we obtain the crystal structure with lattice distortion through structural optimization, which is shown in Fig. 2(e).
@@
After lattice distortion, the symmetry of crystal structure changes from the high-symmetry P4/mmm to the low-symmetry P-421m space group symmetry, corresponding to the D2d point group. Interestingly, the magnetic ground state of monolayer WO with lattice distortion remains d-wave altermagnetic (Fig. 2(e)) (detailed calculations and analysis of d-wave altermagnetism are shown in the supplementary material (SM)). Then, we calculated the phonon spectrum of the supercell WO with altermagnetism, which is shown in Fig. 2(f). From Fig. 2(f), there are no imaginary frequencies in the phonon spectrum. Note: The tiny imaginary frequency at the point is caused by the finite displacement calculation method, and the density functional perturbation theory based on the QE software can completely eliminate the imaginary frequency at (the calculation results are shown in the SM). Therefore, the two imaginary frequencies at the M and points in the normal state of monolayer WO are indeed caused by a CDW. That is to say, monolayer WO is a material in which CDW and altermagnetism coexist. On the other hand, we also calculate the phonon spectrum of the supercell with lattice distortion for nonmagnetic monolayer WO, which is shown in the SM. Our results indicate there are pronounced imaginary frequencies for the supercell with lattice distortion in nonmagnetic monolayer WO, reflecting that nonmagnetic monolayer WO does not host a CDW. Therefore, altermagnetism stabilizes the CDW.
@@
A natural question is what causes the CDW in altermagnetic WO. This requires us to further investigate the electronic structure of altermagnetic WO. In the subsequent electronic-structure calculations, we include a Hubbard U = 2.2 eV to account for the on-site Coulomb repulsion among the 5 orbitals of W atom according to the previous works [53, 54]. We first calculated the electronic band structure and density of states of altermagnetic WO in the normal state, which are shown in Fig. 3(a). From Fig. 3(a), monolayer WO is an altermagnetic semimetal with four pairs of crossing points near the Fermi level. Two pairs of crossing points with opposite polarization are protected by the spin symmetry on the X-M axis, and the other two pairs of crossing points with opposite polarization are protected by the spin symmetry on the Y-M axis. Therefore, monolayer WO is a bipolarized Weyl semimetal, similar to monolayer Fe2WTe4. On the other hand, since the bands near the Fermi level are mainly contributed by the 5 orbitals of W atoms (Fig. 3(a)), monolayer WO is likely to have strong SOC effect. However, calculation of the Lindhard response function shows that monolayer WO exhibits no CDW instability driven by Fermi-surface nesting. In view of the strong phonon softening observed in the normal state, the CDW in monolayer WO is likely induced by strong electron–phonon coupling.
@@
Before calculating the band structure with SOC, we need to first determine the direction of the easy magnetization axis of monolayer WO. This is because the direction of the easy magnetization axis determines the magnetic point group symmetry, which in turn has a significant impact on the electronic structure and topological properties of monolayer WO. Our calculations show that the easy magnetization axis of WO is along the z-axis. So altermagnetic WO breaks the and symmetries under SOC, which in turn leads to the Weyl points near the Fermi level opening gap. Next, we calculated the electronic structure of monolayer WO with SOC, which is shown in Fig. 3(b). Comparing Fig. 3(a) and Fig. 3(b), monolayer WO indeed has strong SOC effect, especially the two pairs of Weyl points near the X and Y points open a band gap of 266 meV. However, monolayer WO remains a semimetal, with its corresponding Fermi surface shown in Fig. 3(c). From Fig. 3(c), the hole Fermi surface is near the point, while the electron Fermi surfaces are located near the X and Y points. Clearly, the hole Fermi surface and the electron Fermi surface cannot be nested by the (, ) wave vector, but the electron Fermi surfaces near the X and Y points can be nested by the (, ) wave vector. Therefore, the CDW in altermagnetic WO is likely caused by Fermi surface nesting.
@@
Typically, CDW degrades the metallic properties of a material, such as transforming a metal into an insulator. Thus, the normal-state altermagnetic semimetal WO should transition to an altermagnetic insulator through the CDW. To verify this speculation, we calculated the electronic structure of the altermagnetic WO without SOC in the CDW state, which is shown in Fig. 3(d). Surprisingly, the altermagnetic WO transitions from the normal-state semimetal phase to a metallic phase in the CDW state, which is quite anomalous. Unlike the normal state, the bands along the -X and -Y directions in the CDW state of WO are spin degenerate, while the bands along the -M direction are spin split. This is because the spin symmetries and protect the spin degeneracy of the bands along the -X and -Y directions in the CDW state of monolayer WO. On the other hand, since the states near the Fermi level are still contributed by the 5 orbitals of W atoms, we also calculated the electronic band structure of the CDW state of WO with SOC (Fig. 3(e)). Although the SOC splits the crossing points along the -M direction and the spin degeneracy on the high-symmetry lines, the splitting is not significant. Additionally, we provided the Fermi surface of the CDW state of WO, which intuitively shows that the CDW state of WO becomes a good metal (Fig. 3(f)).
Another important question is why the CDW in altermagnetic WO leads to the anomalous transition from a semimetal to a metal. To answer this question, we calculated the 5 orbitals of the W atoms projections for bands of the normal state and the CDW state, which are shown in Fig. 4(a) and (b). Comparing the orbital projections of bands of the normal state and the CDW state, we find that the , , and orbitals decrease in energy, while the and orbitals increase in energy in the CDW state. This not only stabilizes the CDW state but also transforms monolayer WO from a semimetal to a metal. Moreover, the structural distortion of WO in the CDW state also supports the decrease in energy of the , , and orbitals and the increase in energy of the and orbitals. Finally, monolayer WO in the CDW state is likely to have strong electron-phonon coupling, and combined with its good metallic properties, this may lead to the coexistence of superconductivity and altermagnetism in the CDW state of WO. That is to say, it may be possible to realize time-reversal symmetry-breaking superconductivity in the CDW state of monolayer WO, which requires further investigation in the future.
Conclusion. Based on symmetry analysis and first-principles calculations, we predict that monolayer WO is a material in which altermagnetism and CDW coexist. Moreover, our calculations reveal that the altermagnetic order in monolayer WO stabilizes the CDW. Further, the CDW is not driven by Fermi-surface nesting but rather by strong electron–phonon coupling. More importantly, the CDW causes monolayer WO to transition from a semimetal to a metal. Combining the strong electron-phonon coupling and good metallic properties in the altermagnetic CDW state, our work may provide new insights into the realization of non-trivial altermagnetic superconductivity.

### v2

# Anomalous charge density wave in altermagnetism

@@ Abstract
Exploring the intricate interplay between magnetism and charge density waves has long been a fundamental pursuit at the forefront of condensed matter research. In this letter, based on symmetry analysis and first-principles calculations, we propose for the first time that anomalous charge density wave can be realized in two-dimensional altermagnetic WO. The anomalous charge density wave is characterized by three key features: (i) Unlike conventional charge density wave, whose stabilization is driven by the opening of a gap near the Fermi level, the anomalous charge density wave is stabilized by the occupied states with energies shifting lower far away from the Fermi level; (ii) the anomalous charge density wave increases the density of states near the Fermi level and then enhances—rather than diminishes—the metallicity of materials; (iii) altermagnetism plays a crucial role in stabilizing anomalous charge density wave. Thus, our work offers a pathway for exploring both the realization and the underlying mechanisms of anomalous charge density waves in magnetic systems.
@@
Introduction. Altermagnetism, a new magnetic phase distinct from both ferromagnetism and antiferromagnetism, has been proposed theoretically and confirmed experimentally [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Altermagnetic materials exhibit momentum-space anisotropic spin splitting without spin–orbit coupling (SOC) and real-space antiparallel spin alignment with zero net total magnetic moment, thereby sharing the advantages of both ferromagnetic and antiferromagnetic materials. Consequently, numerous exotic phenomena can emerge in altermagnetic materials, such as anisotropic spin splitting leading to spin currents [11, 12, 13, 14], giant and tunneling magnetoresistance [15, 16, 17], and piezomagnetic effects [18]. With SOC, due to the break of the time-reversal symmetry, altermagnetic materials can realize the anomalous Hall [19, 20, 21, 22], anomalous magneto-optical [23, 24], anomalous Nernst [25], anomalous thermal Hall [26] and crystal valley Hall [27] effects. On the other hand, the combination of altermagnetism with other matter phases has also attracted widespread interest. The anisotropic spin-splitting altermagnetism without time-reversal symmetry merging with topology gives rise to new topological phases[21, 28, 29, 30, 31, 32, 33]. Coupling time-reversal-broken altermagnetism with superconductivity can lead to nontrivial superconducting states [34, 35]. Altermagnetic ferroelectricity enables electrical control of magnetism [36, 37, 38, 39].
Charge density wave (CDW) is a common quantum phase that accompanies lattice distortion and generally drives a metal-to-insulator transition or a reduction in the density of states (DOS) near the Fermi level (i.e., poorer metallicity) [40, 41, 42, 43, 44], as illustrated in Fig. 1(a). Due to strong electron-phonon coupling, the suppression of the CDW often precedes the emergence of superconductivity [45, 46]. An intriguing possibility arises if, instead of weakening metallicity, the CDW actually enhances metallicity, thereby transforming the traditionally competitive relationship between CDW and superconductivity into a cooperative one. Conventionally, CDW stabilization stems from the opening of a gap near the Fermi level, which lowers the total energy at the cost of reduced metallicity. If the stability of CDW phase is instead driven by the downward shift of locally occupied bands far from the Fermi level, the metallicity of materials may be enhanced, as illustrated in Fig. 1(b). We term such a CDW as an anomalous CDW. In nonmagnetic materials, CDWs typically degrade metallicity; the introduction of magnetism may facilitate the realization of anomalous CDWs. Very recently, the combination of altermagnetism and CDWs has begun to be explored [47, 48, 49]. Thus, whether or not is the anomalous CDW realized in an altermagnetic material?
In this letter, based on symmetry analysis and first-principles calculations, we propose for the first time that anomalous CDW can be realized in two-dimensional altermagentic WO. Moreover, our calculations reveal that altermagnetism plays a crucial role in stabilizing the CDW in monolayer WO. Thus, our work offers a pathway for exploring both the realization and the underlying mechanisms of anomalous charge density waves in magnetic systems.
@@
Results and analysis. Recently, monolayer CrO and MoO have been predicted to be bipolarized Weyl semimetals with altermagnetism, and they will transition to the Chern insulator phase under strain [50, 21, 51, 29]. Monolayer CrO is a perfect bipolarized Weyl semimetal, while the overlap of the valence and conduction bands in monolayer MoO leads to the presence of both electron and hole Fermi surfaces. Therefore, monolayer MoO exhibits better metallic properties than monolayer CrO. If W atoms substitute Mo atoms, monolayer WO should have stronger metallic properties. Additionally, 5 transition metal oxides may lead to structural distortions [52], so monolayer WO provides the possibility for the coexistence of altermagnetism and anomalous CDW.
@@
Similar to the monolayer CrO, the crystal structure of monolayer WO has the space group symmetry, corresponding to the point group. The primitive cell of monolayer WO contains two W atoms and two O atoms, with an optimized lattice constant of 3.696 Å. Moreover, the detailed calculations reveal that the magnetic ground state of monolayer WO remains -wave altermagnetic across a range of Hubbard U, consistent with monolayer CrO (the results and analyses are provided in the Supplementary Material (SM) [53]). Then, we investigate the dynamical stability of monolayer WO based on altermagnetic structure. We calculate the phonon spectrum of WO, as shown in Fig. 2(a). From Fig. 2(a), WO has two phonon modes with imaginary frequencies respectively at the high-symmetry M and points, and the imaginary frequency at the M point is larger than that at the point. We speculate that the imaginary frequencies might be caused by a CDW. To verify this speculation, we first calculate the phonon spectrum of the supercell. Under the supercell, the M point of the primitive cell Brillouin zone (BZ) is folded to the point, as shown in Fig. 2(b). Therefore, the vibration mode with imaginary frequency at the M point of the primitive cell BZ is folded to the point. Then, we plot these two vibration modes with imaginary frequency, as shown in Fig. 2(c) and 2(d), respectively. From Fig. 2(c) and 2(d), these two vibration modes originate from the vibrations of W atoms. The vibration mode with imaginary frequency at the point is the relative vibration of W atoms along the direction, and the vibration mode with imaginary frequency at the M point is the in-plane relative vibration of W atoms. Based on these two vibration modes, we obtain the crystal structure with lattice distortion through structural optimization, as shown in Fig. 2(e).
@@
After lattice distortion, the symmetry of crystal structure changes from the high-symmetry to the low-symmetry - space group symmetry, corresponding to the point group. Interestingly, the magnetic ground state of monolayer WO with lattice distortion remains -wave altermagnetic (Fig. 2(e)) (the results and analyses are provided in the SM [53]). Then, we calculate the phonon spectrum for this distorted supercell with altermagnetism, as shown in Fig. 2(f). From Fig. 2(f), there are no imaginary frequencies in the phonon spectrum. Note: The tiny imaginary frequency at the point is caused by the finite displacement calculation method, and the density-functional perturbation theory based on the Quantum Espresso [54] can completely eliminate the imaginary frequency at point (the results and analyses are provided in the SM [53]). Therefore, the two imaginary frequencies at the M and points in the normal phase of monolayer WO are indeed caused by a CDW. That is to say, monolayer WO is a material in which CDW and altermagnetism coexist. On the other hand, we also calculated the phonon spectrum of nonmagnetic monolayer WO in the CDW phase, as shown in Fig. 2(g). From Fig. 2(g), there are pronounced imaginary frequencies, reflecting that nonmagnetic monolayer WO does not host a CDW. Therefore, altermagnetism plays a crucial role in stabilizing the CDW in monolayer WO.
@@
A natural question is what causes the CDW in altermagnetic WO. This requires us to further investigate the electronic structure of altermagnetic WO. In the subsequent electronic structure calculations, we include a Hubbard U = 2.2 eV to account for the on-site Coulomb repulsion among the 5 orbitals of W atoms according to previous studies [55, 56]. We first calculate the electronic band structure and density of states of altermagnetic WO without SOC in the normal phase, as shown in Fig. 3(a). From Fig. 3(a), monolayer WO is an altermagnetic semimetal with four pairs of crossing points near the Fermi level. Two pairs of crossing points with opposite polarization are protected by the spin symmetry along the X-M direction, and the other two pairs of crossing points with opposite polarization are protected by the spin symmetry along the Y-M direction. Therefore, monolayer WO is a bipolarized Weyl semimetal, similar to monolayer Fe2WTe4 [29]. On the other hand, since the bands near the Fermi level are mainly contributed by the 5 orbitals of W atoms (Fig. 3(a)), monolayer WO likely exhibits strong SOC effect.
@@
Before calculating the band structure with SOC, we need to first determine the direction of the magnetic easy axis of monolayer WO. This is because the direction of the magnetic easy axis determines the magnetic point group symmetry, which in turn has a significant impact on the electronic structure and topological properties of monolayer WO. Our calculations show that the magnetic easy axis of WO is along the direction. So altermagnetic WO breaks the and symmetries under SOC, which in turn leads to the Weyl points near the Fermi level opening gap. Next, we calculate the electronic structure of monolayer WO with SOC, as shown in Fig. 3(b). Comparing Fig. 3(a) and 3(b), monolayer WO indeed exhibits strong SOC effect, especially the two pairs of Weyl points near the X and Y points open a band gap of 266 meV. However, monolayer WO remains a semimetal, with its corresponding Fermi surface shown in Fig. 3(c). From Fig. 3(c), the hole Fermi surface is near the point, while the electron Fermi surfaces are located near the X and Y points. Clearly, the hole Fermi surface and the electron Fermi surface cannot be nested by the (, ) wave vector, but the electron Fermi surfaces near the X and Y points can be nested by the (, ) wave vector. However, calculation of the Lindhard susceptibility function shows that monolayer WO exhibits no CDW instability driven by Fermi surface nesting (the results and analyses are provided in the SM [53]). In view of the strong phonon softening observed in the normal phase, the CDW in monolayer WO likely arises from the synergy of electron-phonon coupling with altermagnetism.
@@
Is the CDW in monolayer WO an anomalous CDW? To verify it, we calculate the electronic structure of the altermagnetic WO without SOC in the CDW phase, as shown in Fig. 3(d). Unlike the normal phase, the bands along the -X and -Y directions in the CDW phase of WO are spin degenerate, while the bands along the -M direction are spin split (Fig. 3(d)). This is because the spin symmetries and protect the spin degeneracy of the bands along the -X and -Y directions in the CDW phase of monolayer WO. On the other hand, since the states near the Fermi level are still contributed by the 5 orbitals of W atoms, we also calculate the electronic band structure of the CDW phase of WO with SOC (Fig. 3(e)). Although the SOC splits the crossing points along the -M direction and the spin degeneracy on the high-symmetry lines, the splitting is not significant. Importantly, the DOS in the CDW phase is 2.32 times that of the normal phase at the Fermi level, transforming monolayer WO from a semimetal into a metal. Moreover, we present the Fermi surface of the CDW phase of WO, which intuitively shows that the CDW phase of WO becomes a good metal (Fig. 3(f)). Therefore, the anomalous CDW can be realized in altermagnetic WO. In addition, we also calculate the electronic structure of nonmagnetic monolayer WO in the CDW phase (the results and analyses are provided in the SM [53]). In the nonmagnetic phase, monolayer WO remains metallic; its bands differ markedly from those in the altermagnetic phase, yet these states near the Fermi level are still dominated by the 5 orbitals of W atoms. Despite these similarities, the CDW is unstable in the nonmagnetic case. This further underscores the crucial role played by altermagnetism in enabling the anomalous CDW.
Another important question is why the CDW in altermagnetic WO is anomalous. To answer this question, we first compare the crystal-field splittings of the 5 orbitals from W atoms before and after the lattice distortion. In the normal-phase monolayer WO, W sits in a rectangular crystal field whose and orbitals lie at higher energy, while , , and are lower (Fig. 4(a)). After the CDW distortion, W sits in a distorted rectangular field that pushes and even higher and pulls , , and to lower energies (Fig. 4(a)). Moreover, the large downward shift stabilizes the CDW phase. Then, we plot band structures with -orbital projection for normal-phase and CDW-phase WO, as shown in Fig. 4(b) and 4(c), respectively. They confirm the higher-lying and and lower-lying , , and levels, and show that after the CDW transition and rise further while , , and down further (Fig. 4(b) and 4(c)). Notably, the bands crossing the Fermi level are dominated by the and orbitals (Fig. 4(b) and 4(c)); their upward shift in the CDW phase enlarges the Fermi surface and turns WO into a metal. Therefore, the origin of the anomalous CDW in monolayer WO is fully consistent with the proposed mechanism illustrated in the schematic (Fig. 1(b)).
According to the results and analysis above, the anomalous CDW in monolayer WO exhibits three distinctive features: (i) Altermagnetism plays a crucial role in stabilizing anomalous CDW; (ii) Unlike conventional CDW, whose stabilization is driven by the opening of a gap near the Fermi level, the anomalous CDW is stabilized by the occupied states with energies shifting lower far away from the Fermi level; and (iii) The second feature increases the density of states near the Fermi level, thereby driving the anomalous semimetal-to-metal transition in monolayer WO induced by CDW. Furthermore, the three distinctive features of the anomalous CDW in monolayer WO are not confined to altermagnetic materials; they may also extend to other magnetic phases, such as collinear antiferromagnetism and unconventional non-collinear antiferromagnetism. Our work therefore offers a promising avenue for exploring anomalous charge density waves in magnetic systems more broadly.